### v1

@@ 4.2 ADMM solver for the subproblem in DCEN-DCA
(30) $\bm{x}_{k+1}\in\arg\min_{\bm{x}\in\mathbb{R}^{n}}\;\frac{1}{2}\|\bm{Ax}-\bm{b}\|_{2}^{2}+\lambda\left(\gamma\|\bm{x}\|_{1}+\frac{3(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}\right)-\lambda\left\langle\bm{x},d_{\ell_{2}}(\bm{x}_{k})\right\rangle.$
@@
$\min_{\bm{x},\bm{z}\in\mathbb{R}^{n}}\;\frac{1}{2}\|\bm{Ax}-\bm{b}\|_{2}^{2}+\frac{3\lambda(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}-\lambda\langle\bm{x},\bm{d}_{\ell_{2}}(\bm{x}_{k})\rangle+\lambda\gamma\|\bm{z}\|_{1}\;\;\text{s.t.}\;\bm{x}-\bm{z}=\bm{0}.$
The augmented Lagrangian function with penalty parameter $\rho>0$ and dual variable $\bm{y}$ is given by $\mathcal{L}_{\rho}(\bm{x},\bm{z},\bm{y})=\frac{1}{2}\|\bm{Ax}-\bm{b}\|_{2}^{2}+\frac{3\lambda(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}-\lambda\langle\bm{x},\bm{d}_{\ell_{2}}(\bm{x}_{k})\rangle+\lambda\gamma\|\bm{z}\|_{1}+\langle\bm{y},\bm{x}-\bm{z}\rangle+\frac{\rho}{2}\|\bm{x}-\bm{z}\|_{2}^{2}$. Let $f(\bm{x})=\frac{1}{2}\|\bm{Ax}-\bm{b}\|_{2}^{2}+\frac{3\lambda(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}$. Then at the $t$-th iteration of ADMM (with the outer DCEN-DCA index $k$ suppressed for clarity), the primal and dual variables are updated according to
(31) $\begin{cases}\bm{x}_{t+1}=\displaystyle\arg\min_{\bm{x}\in\mathbb{R}^{n}}\;f(\bm{x})-\lambda\bm{x}^{\top}\bm{d}_{\ell_{2}}(\bm{x}_{k})+\bm{y}_{k}^{\top}\bm{x}+\frac{\rho}{2}\|\bm{x}-\bm{z}_{k}\|_{2}^{2},\\ \bm{z}_{t+1}=\displaystyle\arg\min_{\bm{z}\in\mathbb{R}^{n}}\;\frac{\rho}{2}\|\bm{x}_{k+1}-\bm{z}\|_{2}^{2}+\lambda\gamma\|\bm{z}\|_{1}+\langle\bm{y}_{k},\bm{x}_{k+1}-\bm{z}\rangle,\\ \bm{y}_{t+1}=\bm{y}_{t}+\rho\left(\bm{x}_{t+1}-\bm{z}_{t+1}\right).\end{cases}$
As for the $\bm{x}$-subproblem of (31), the closed form solution is $\bm{x}_{t+1}=\big(\bm{A}^{\top}\bm{A}+(3\lambda(1-\gamma)+\rho)\bm{I}\big)^{-1}\left(\bm{A}^{\top}\bm{b}+\lambda\,\bm{d}_{\ell_{2}}(\bm{x}_{k})-\bm{y}_{t}+\rho\bm{z}_{t})\right)$. Let $\eta=3\lambda(1-\gamma)+\rho$ and $\bm{R}_{t}=\bm{A}^{\top}\bm{b}+\lambda\,\bm{d}_{\ell_{2}}(\bm{x}_{k})-\bm{y}_{t}+\rho\bm{z}_{t}$. By applying the Sherman–Morrison–Woodbury (SMW) identity, we have $\bm{x}_{t+1}=\eta^{-1}\bm{R}_{t}-\eta^{-2}\bm{A}^{\top}\big(\bm{I}+\eta^{-1}\bm{A}\bm{A}^{\top}\big)^{-1}\bm{A}\bm{R}_{t}$. For moderate $n$ precompute a Cholesky factorization of the matrix $\bm{A}^{\top}\bm{A}+(3\lambda(1-\gamma)+\rho)\bm{I}$ and for large-scale problems use conjugate gradient (CG) with warm start.
For the $\bm{z}$-update
$\bm{z}_{t+1}$ $=\arg\min_{\bm{z}\in\mathbb{R}^{n}}\;\frac{\rho}{2}\|\bm{x}_{t+1}-\bm{z}\|_{2}^{2}+\lambda\gamma\|\bm{z}\|_{1}+\langle\bm{y}_{t},\bm{x}_{t+1}-\bm{z}\rangle$
$=\arg\min_{\bm{z}\in\mathbb{R}^{n}}\;\|\bm{z}\|_{1}+\frac{\rho}{2\lambda\gamma}\|\bm{z}-(\bm{x}_{t+1}+1/\rho\bm{y}_{t})\|_{2}^{2},$
which has a solution $\bm{z}_{t+1}:=\operatorname{sign}\!\left(\bm{x}_{t+1}+\rho^{-1}\bm{y}_{t}\right)\odot\max\!\left(\left|\bm{x}_{t+1}+\rho^{-1}\bm{y}_{t}\right|-\tfrac{\lambda\gamma}{\rho},\,\bm{0}\right)$.
@@
(32) $\|\bm{r}_{t}\|_{2}\leq\varepsilon_{\mathrm{pri}}^{t},\quad\|\bm{s}_{t}\|_{2}\leq\varepsilon_{\mathrm{dual}}^{t},\;\text{and}\;\;\mathrm{relerr}_{t}<\varepsilon_{\mathrm{rel}},$
@@
(33a) $\varepsilon_{\mathrm{pri}}^{t}=\sqrt{n}\,\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\big(\|\bm{x}_{t}\|_{2},\|\bm{z}_{t}\|_{2}\big),\;\varepsilon_{\mathrm{dual}}^{t}=\sqrt{n}\,\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\|\bm{y}_{t}\|_{2},$
(33b) $\mathrm{relerr}_{t}=\frac{\|\bm{x}_{t}-\bm{z}_{t}\|_{2}}{\max\big(\|\bm{x}_{t}\|_{2},\|\bm{z}_{t}\|_{2},\varepsilon\big)},$
where $\varepsilon_{\mathrm{abs}}$ and $\varepsilon_{\mathrm{rel}}$ denote the absolute and relative tolerances, respectively, and $\varepsilon$ is the machine precision. The above stopping rule follows the standard criterion proposed in [boyd2011distributed]. ADMM for solving the DCEN-DCA subproblem is summarized in Algorithm 2. Since the subproblem is convex, ADMM converges to its global minimizer under standard assumptions; when embedded in DCEN-DCA (outer nonconvex loop) the whole scheme typically converges to a stationary point of the original nonconvex problem under usual DCEN-DCA assumptions [tao1998dc, le2018dc].
Algorithm 2 ADMM for minimizing DCEN-DCA subproblem (30).

### v2

@@ V-B ADMM solver for the subproblem in DCEN-DCA
$\bm{x}_{k+1}\in\arg\min_{\bm{x}\in\mathbb{R}^{n}}g(\bm{x})-\lambda\left\langle\bm{x},d_{\ell_{2}}(\bm{x}_{k})\right\rangle.$ (35)
@@
$\min_{\bm{x},\bm{z}\in\mathbb{R}^{n}}\;\ell(\bm{x})+\frac{3\lambda(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}-\lambda\langle\bm{x},\bm{d}_{\ell_{2}}(\bm{x}_{k})\rangle+\lambda\gamma\|\bm{z}\|_{1}$
$\;\;\text{s.t.}\;\bm{x}-\bm{z}=\bm{0}.$
The augmented Lagrangian function with penalty parameter $\rho>0$ and dual variable $\bm{y}$ is given by $\mathcal{L}_{\rho}(\bm{x},\bm{z},\bm{y})=\ell(\bm{x})+\frac{3\lambda(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}-\lambda\langle\bm{x},\bm{d}_{\ell_{2}}(\bm{x}_{k})\rangle+\lambda\gamma\|\bm{z}\|_{1}+\langle\bm{y},\bm{x}-\bm{z}\rangle+\frac{\rho}{2}\|\bm{x}-\bm{z}\|_{2}^{2}$. Let $f(\bm{x})=\ell(\bm{x})+\frac{3\lambda(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}$. Then at the $t$-th iteration of ADMM (with the outer DCEN-DCA index $k$ suppressed for clarity), the primal and dual variables are updated according to
$\begin{cases}\bm{x}_{t+1}=\displaystyle\arg\min_{\bm{x}\in\mathbb{R}^{n}}\;f(\bm{x})-\lambda\bm{x}^{\top}\bm{d}_{\ell_{2}}(\bm{x}_{k})+\bm{y}_{k}^{\top}\bm{x}+\frac{\rho}{2}\|\bm{x}-\bm{z}_{k}\|_{2}^{2},\\ \bm{z}_{t+1}=\displaystyle\arg\min_{\bm{z}\in\mathbb{R}^{n}}\;\frac{\rho}{2}\|\bm{x}_{k+1}-\bm{z}\|_{2}^{2}+\lambda\gamma\|\bm{z}\|_{1}+\langle\bm{y}_{k},\bm{x}_{k+1}-\bm{z}\rangle,\\ \bm{y}_{t+1}=\bm{y}_{t}+\rho\left(\bm{x}_{t+1}-\bm{z}_{t+1}\right).\end{cases}$ (36)
As for the $\bm{x}$-subproblem of (36), the closed form solution is $\bm{x}_{t+1}=\big(\bm{A}^{\top}\bm{A}+(3\lambda(1-\gamma)+\rho)\bm{I}\big)^{-1}\left(\bm{A}^{\top}\bm{b}+\lambda\,\bm{d}_{\ell_{2}}(\bm{x}_{k})-\bm{y}_{t}+\rho\bm{z}_{t})\right)$. Let $\eta=3\lambda(1-\gamma)+\rho$ and $\bm{R}_{t}=\bm{A}^{\top}\bm{b}+\lambda\,\bm{d}_{\ell_{2}}(\bm{x}_{k})-\bm{y}_{t}+\rho\bm{z}_{t}$. By applying the Sherman–Morrison–Woodbury (SMW) identity, we have $\bm{x}_{t+1}=\eta^{-1}\bm{R}_{t}-\eta^{-2}\bm{A}^{\top}\big(\bm{I}+\eta^{-1}\bm{A}\bm{A}^{\top}\big)^{-1}\bm{A}\bm{R}_{t}$. For moderate $n$ precompute a Cholesky factorization of the matrix $\bm{A}^{\top}\bm{A}+(3\lambda(1-\gamma)+\rho)\bm{I}$ and for large-scale problems use conjugate gradient (CG) with warm start [31].
For the $\bm{z}$-update $\bm{z}_{t+1}=\displaystyle\arg\min_{\bm{z}\in\mathbb{R}^{n}}\;\|\bm{z}\|_{1}+\frac{\rho}{2\lambda\gamma}\|\bm{z}-(\bm{x}_{t+1}+1/\rho\bm{y}_{t})\|_{2}^{2}$, which has a solution $\bm{z}_{t+1}:=\operatorname{sign}(\bm{x}_{t+1}+\rho^{-1}\bm{y}_{t})\odot\max(\left|\bm{x}_{t+1}+\rho^{-1}\bm{y}_{t}\right|-\tfrac{\lambda\gamma}{\rho},\,\bm{0})$.
@@
$\|\bm{r}_{t}\|_{2}\leq\varepsilon_{\mathrm{pri}}^{t},\quad\|\bm{s}_{t}\|_{2}\leq\varepsilon_{\mathrm{dual}}^{t},\;\text{and}\;\;\mathrm{relerr}_{t}<\varepsilon_{\mathrm{rel}},$ (37)
@@
$\varepsilon_{\mathrm{pri}}^{t}=\sqrt{n}\,\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\big(\|\bm{x}_{t}\|_{2},\|\bm{z}_{t}\|_{2}\big),$ (38a)
$\varepsilon_{\mathrm{dual}}^{t}=\sqrt{n}\,\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\|\bm{y}_{t}\|_{2},$ (38b)
$\mathrm{relerr}_{t}=\frac{\|\bm{x}_{t}-\bm{z}_{t}\|_{2}}{\max\big(\|\bm{x}_{t}\|_{2},\|\bm{z}_{t}\|_{2},\varepsilon\big)},$ (38c)
where $\varepsilon_{\mathrm{abs}}$ and $\varepsilon_{\mathrm{rel}}$ denote the absolute and relative tolerances, respectively, and $\varepsilon$ is the machine precision. The above stopping rule follows the standard criterion proposed in [3]. ADMM for solving the DCEN-DCA subproblem is summarized in Algorithm 2. Since the subproblem is convex, ADMM converges to its global minimizer under standard assumptions; when embedded in DCEN-DCA (outer nonconvex loop) the whole scheme typically converges to a stationary point of the original nonconvex problem under usual DCEN-DCA assumptions [32, 21].
Algorithm 2 ADMM for DCEN-DCA subproblem (35).